15 Campbell Manufacturing started in 2018 with the following account balances: Cash $5,700 5,120 6, 000 2,200 Common stock Retained earnings Raw materials inventory Work in process inventory Finished goods inventory (400 units $5.90 each) 860 2,360 Transactions during 2018 1. Purchased $2,890 of raw materials with cash. 2. Transferred $3,770 of raw materials to the production department. 3. Incurred and paid cash for 180 hours of direct labor$15.20 per hour 4. Applied overhead costs to the Work in Process Inventory account. The predetermined overhead rate is $16.40 per direct labor hour 5. Incurred actual overhead costs of $3,000 cash. 6. Completed work on 1,280 units for $5.50 per unit. 7. Paid $1,190 in selling and administrative expenses in caslh 8. Sold 1,280 units for $10,400 cash revenue (assume FIFO cost flow). 9. Campbell charges overapplied or underapplied overhead directly to Cost of Goods Sold. Required a. Record the preceding events in a horizontal statements model. The beginning balances are shown as an example b. Prepare a schedule of cost of goods manufactured and sold, an income statement, and a balance sheet for 2018 Complete this question by entering your answers in the tabs below.
15 Complete this question by entering your answers in the tabs below Req B Inc Req B Bal Sheet Req B CGM Sched Req A Stmt Record the preceding events in a horizontal statements model. The beginning balances are shown as an example. (Enter decreases to account bal Equity Asscts Event No Raw Work in Finished Common Retained Net Cash + Manufacturing Revenue Expense Materials Process Goods Stock Earnings B.B. 5,700 +S 2,200S 860 2.360$ 5,1206,000 2. 3. 5. 6. 8b Total Req A Req B CGM Sched >
Req B CGMReq B Inc Req B Bal Sched Sheet Prepare a schedule of cost of goods manufactured and sold for 2018. CAMPBELL MANUFACTURING COMPANY Cost of Goods Manufactured and Sold For the Year Ended December, 2018 Raw materials available Raw materials used Total manufacturing costs Total work in process inventory Cost of goods manufactured Cost of goods available for sale Cost of goods sold Req B CGM Sched
ED 15 Req B CGM Sched Rq IncReq B Bal Stmt Req A Sheet Prepare a balance sheet for 2018. CAMPBELL MANUFACTURING COMPA Balance Shect At December, 2018 Assets Total assets Equity Total equity Req B Inc Stmt Req B Bal Sheet
Answer #1

a.

Assets = Equity
Event No. Cash Manufacturing
Overhead
Raw
Materials
Work in
Process
Finished
Goods
= Common Stock Retained Earnings Revenue Expense Net Income
B.B. 5700 2200 860 2360 = 5120 6000
1 -2890 2890 =
2 -3770 3770 =
3 -2736 2736 =
4 -2952 2952 =
5 -3000 3000 =
6 -7040 7040 =
7 -1190 = -1190 1190 -1190
8a. 10400 = 10400 10400 10400
8b. -7200 = -7200 7200 -7200
9 -48 = -48 48 -48
Total 6284 0 1320 3278 2200 = 5120 7962 10400 8438 1962

Cost of goods sold = (400 x $5.90) + (880 x $5.50) = $2360 + $4840 = $7200

b.

CAMPBELL MANUFACTURING COMPANY
Cost of Goods Manufactured and Sold
For the Year Ended December, 2018
Beginning raw materials invenotry 2200
Add purchases 2890
Raw materials available 5090
Less ending balance of raw materials 1320
Raw materials used 3770
Direct labor 2736
Manufactuirng overhead 2952
Total manufacturing costs 9458
Add beginning work in process inventory 860
Total work in process inventory 10318
Less ending work in process inventory 3278
Cost of goods manufactured 7040
Add beginning finished goods inventory 2360
Cost of goods available for sale 9400
Less ending finished goods inventory 2200
Cost of goods sold 7200
CAMPBELL MANUFACTURING COMPANY
Income Statement
For the Year Ended December, 2018
Sales revenue 10400
Cost of goods sold 7248
Gross profit 3152
Selling and administrative expenses 1190
Net income 1962
CAMPBELL MANUFACTURING COMPANY
Balance Sheet
At December, 2018
Assets
Cash 6284
Raw materials inventory 1320
Work in process inventory 3278
Finished goods inventory 2200
Total assets 13082
Equity
Common stock 5120
Retained earnings 7962
Total equity 13082

Note: All amounts have been shown as positive amounts in the absence of specific instructions regarding the same. Kindly fill in deductions with a minus sign if required.